8. The Oriole Acres Inn is trying to determine its break-even point during its off-peak season. The inn has 50 rooms that it rents at $44 a night. Operating costs are as follows:

Salaries $5,400 per month
Utilities $1,000 per month
Depreciation $1,200 per month
Maintenance $1,288 per month
Maid service $7 per room
Other costs $15 per room

Determine the inn’s break-even point in number of rented rooms per month.

Break-even point

enter the break-even point in number of rented rooms

rooms

  

  

Determine the inn’s break-even point in dollars.

Break-even point in

$enter the break-even point in dollars
Answer #1

Contribution Margin per Room per Night = Room Rent per Night per Room - Variable Cost per Room per Night

Room Rent per Night per Room = $ 44

Variable Costs per Room = Maid Service Cost per Room + Other Cost per Room

Variable Costs per Room = 7 + 15 = $ 22

Contribution Margin per Night per Room = 44 - 22 = $ 22

Total Fixed Costs = Salaries + Utilities + Depreciation + Maintenance

Total Fixed Costs = 5,400 + 1,000 + 1,200 + 1,288 = $ 8,888

Break Even Point in Number of Rented Rooms = Total Fixed Costs / Contribution Margin per Room

Break Even Point in Number of Rented Rooms = 8,888 / 22 = 404 Rooms

Break Even Point in Dollars = Total Fixed Costs / Contribution Margin Ratio

Break Even Point in Dollars = 8,888 / 50%

Break Even Point in Dollars = $ 17,776

Contribution Margin Ratio = Contribution Margin / Room Rent * 100

Contribution Margin Ratio = 22 / 44 * 100

Contribution Margin Ratio = 50%